# Common Envelope Evolution of Ultralight Boson Clouds

### v1

@@ Appendix A Eigenstates of Gravitational Molecules
with $a$ being the semi-major axis. The Laplacian in this coordinate can be written as
@@
for $\bar{\psi}_{n\ell m}$ in Eq. (11), Eq. (10) becomes separable, and leads to
@@ A.1 Equal-mass System
For equal-mass binaries, the angular part $S_{\ell m}(\zeta)$ turns out to be the spheroidal wave function, with a certain correspondence between $A_{\ell m}$ and $\tilde{\epsilon}$. For the radial part, the asymptotic behavior of $R_{n\ell}(\xi)$ should converge to 0 as $\xi\rightarrow\infty$, and the number of nodes is nl1n-l-1. By imposing the boundary condition in the limit of $\xi\rightarrow\infty$, we can determine the value of eigenenergy $\tilde{\epsilon}$ and hence $A_{\ell m}$, and also can solve the wavefunction $\bar{\psi}_{n\ell m}$ numerically, up to a normalization which can be fixed by requiring
@@
The parity of the gravitational molecules depends on the angular part. To be specific,
$S_{\ell m}(\zeta)$ $=S_{\ell m}(-\zeta),\quad\ell+m\ \rm{is\ even;}$ (51)
$S_{\ell m}(\zeta)$ $=-S_{\ell m}(-\zeta),\quad\ell+m\ \rm{is\ odd.}$
This property, along with the exponential term of $\chi$ indicates the selection rules for the coupling between molecule states shown in Eq. (18).
@@ A.2 General Mass-ratio System
For cases with $q\neq 1$, we can further convert Eq. (49) into standard forms by introducing
@@
$\lambda=\xi+1,\quad G(\lambda)=\left(\lambda^{2}-2\lambda\right)^{-m/2}R(\lambda-1).$
Following Ref. [47], $F(\kappa)$ and $G(\lambda)$ are expanded as
@@
By plugging Eq. (52) and Eq. (53) into Eq. (49), we obtain the recurrence relation
$0=$ $D^{i}_{0}-\frac{C^{i}_{0}E^{i}_{1}}{D^{i}_{1}-}\frac{C^{i}_{1}E^{i}_{2}}{C^{i}_{2}-}\frac{C^{i}_{2}E^{i}_{3}}{D^{i}_{3}-}$.\,, (54)
@@
By solving Eq. (54) numerically, we can obtain the energy and angular eigenvalues for gravitational molecules at a certain orbital separation.

### v2

@@ Appendix A EIGENSTATES OF GRAVITATIONAL MOLECULES
with $a$ being the semimajor axis. The Laplacian in this coordinate can be written as
@@
for $\bar{\psi}_{n\ell m}$ in Eqs. (11), (10) becomes separable and leads to
@@ A.1 Equal-mass system
For equal-mass binaries, the angular part of $\psi_{n\ell m}$, $S_{\ell m}(\zeta)$, turns out to be the spheroidal harmonics. To obtain the radial function $R_{n\ell}(\xi)$, we first introduce
$\tilde{R}(\xi)=(\xi^{2}-1)^{-|m|/2}R(\xi).$ (50)
In the limit of $\xi\rightarrow 1$, Eq. (49) leads to
$\partial_{\xi}\tilde{R}=-\frac{2\tilde{a}(1+q)+\tilde{a}^{2}\tilde{\epsilon}/2+|m|(|m|+1)-A_{\ell m}}{2(|m|+1)}\tilde{R}$. (51)
As $\xi\rightarrow\infty$, we expect $\tilde{R}$ approaches as we look for bound states. Then the radial function $R$ and the eigenenergy of the molecular states can be obtained by solving Eq. (49) with the shooting method: We integrate Eq. (49) starting from $\xi_{i}$, which is very close to $1$, with the boundary condition derived from (51). $\tilde{R}(\xi_{i})$ could be chosen as any constant, and will be fixed later with the normalization condition
@@
Then by fine-tuning $\tilde{\epsilon}$, we can find a solution that approaches 0 at large $\xi$ (see Fig. 8 for illustration). Here $A_{\ell m}$ is determined by $\tilde{\epsilon}$ given the spheroidal harmonics.
Figure 8: The shooting method used to determine the eigenenergy of the gravitational molecule. The vertical axis is the value of the radial part $R_{21}(\xi)$ of the molecular state $\bar{\Psi}_{21\pm 1}$. The dashed, solid, and dotted lines correspond to $\tilde{\epsilon}=-0.702149$, 0.701638-0.701638, and 0.70119-0.70119, and the middle one is the best value for its convergence.
@@ A.2 General mass-ratio system
For $q\neq 1$, the eigenvalues are obtained with the continued fraction method. Following Ref. [56], we introduce
@@
$\lambda=\xi+1,\quad G(\lambda)=\left(\lambda^{2}-2\lambda\right)^{-m/2}R(\lambda-1)\,$,
and expand $F(\kappa)$ and $G(\lambda)$ as
@@
By plugging Eqs. (53) and  (54) into Eq. (49), we obtain the recurrence relation
$0=$ $D^{i}_{0}-\frac{C^{i}_{0}E^{i}_{1}}{D^{i}_{1}-}\frac{C^{i}_{1}E^{i}_{2}}{C^{i}_{2}-}\frac{C^{i}_{2}E^{i}_{3}}{D^{i}_{3}-}\ldots\,$, (55)
@@
The roots of Eq. (55) can be obtained starting from the limit of $\tilde{a}\ll 1$, where we have $\tilde{\epsilon}\simeq\alpha^{\prime 2}\mu/(2n^{2})$ with $\alpha^{\prime}=GM_{1}(1+q)\mu$. Then by stepping up $\tilde{a}$ in small increments, we can obtain the splitting structure of energy and angular eigenvalues for gravitational molecules at a certain orbital separation.
A.3 Convergence test
For both shooting and continued fraction methods, the convergence can be tested with
$e_{n}\equiv|\tilde{\epsilon}_{n}-\tilde{\epsilon}_{n-1}|\,.$ (57)
For the shooting method, $n$ represents the iterations of fine-tuning the shooting parameters. For the continued fraction method, $n$ is the cutoff order of Eq. (55). When computing the eigenenergies, we demand $e_{n+1}/\tilde{\epsilon}_{n}\lesssim 10^{-2}$ to guarantee the precision. In addition, we examined
$p(n)\equiv\frac{\log|e_{n+1}-e_{n}|}{\log|e_{n}-e_{n-1}|}\,$, (58)
which approaches $1$ at large $n$, indicating the linear convergency of the results [64]. For example, in Figs. 9 and  10, we show the behavior of $e_{n}$ and $p_{n}$ for certain numerical results. The other results manifest a similar convergency behavior.
Figure 9: Convergence test of the shooting method used to solve the eigenenergy of $\bar{\psi}_{210}$ at the separation $\tilde{a}=0.1$. Both $e_{n+1}/\tilde{\epsilon}_{n}$ approaching zero and $p_{n}$ approaching 11 indicate the convergency of the numerical computation.
Figure 10: Convergence test of the continued fraction method used to solve the eigenenergy of $\bar{\psi}_{211}$ for $q=1$. The $e_{n+1}/\tilde{\epsilon}_{n}$ approaches zero indicating the convergency of the numerical computation, where colors of curves represent the eigenenergy at various orbital separations
Appendix B GRAVITATIONAL ATOM IN PARABOLIC COORDINATE
While the separable eigenfunctions of the gravitational atom are usually studied in spherical coordinates, they can also be obtained in parabolic coordinates [54] . The parabolic coordinates are defined by
$x=\sqrt{\rho\sigma}\cos{\chi},\quad y=\sqrt{\rho\sigma}\sin{\chi},\quad z=\frac{1}{2}(\rho-\sigma),$ (59)
and the time-independent Schrödinger equation in the coordinates can be written as
$E\bar{\varphi}=$ $-\frac{2}{\mu(\rho+\sigma)}\left[\frac{\partial}{\partial\rho}\left(\rho\frac{\partial}{\partial\rho}\right)+\frac{\partial}{\partial\sigma}\left(\sigma\frac{\partial}{\partial\sigma}\right)\right]\bar{\varphi}$ (60)
$-\frac{1}{2\mu\rho\sigma}\frac{\partial^{2}}{\partial\chi^{2}}\bar{\varphi}-\frac{2\alpha}{\rho+\sigma}\bar{\varphi}.$
By substituting the ansatz
$\bar{\varphi}=R_{p}(\rho)S_{p}(\sigma)e^{im\chi}$ (61)
into Eq. (60), we can separate the variables and get
$\frac{d}{d\rho}\left(\rho\frac{dR_{p}}{d\rho}\right)+\left(\frac{E\mu}{2}\rho-\frac{m^{2}}{4\rho}+\beta_{1}\right)R_{p}$ $=0,$ (62)
$\frac{d}{d\sigma}\left(\sigma\frac{dS_{p}}{d\sigma}\right)+\left(\frac{E\mu}{2}\sigma-\frac{m^{2}}{4\sigma}+\beta_{2}\right)S_{p}$ $=0,$
where $\beta_{1}+\beta_{2}=\alpha\mu$. The solutions to Eq. (62) is known [65], and the eigenstates are given by
$\bar{\varphi}_{n_{s}km}$ $=e^{im\chi}(\rho\sigma)^{|m|/2}L^{|m|}_{k}(\frac{\sigma}{n})L^{|m|}_{n_{s}-k-|m|-1}(\frac{\rho}{n})e^{-(\rho+\sigma)/(2n_{s})},$ (63)
with $E=-\frac{\alpha^{2}\mu}{2n_{s}^{2}}$, where $L$ is the associated Laguerre polynomials. The relation between the eigenstates (63) and the usual eigenstates obtained in the spherical coordinates, i.e., $\bar{\varphi}_{nlm}$, has been investigated in Ref. [54]. Here we simply present the results
$\bar{\varphi}_{nlm}=\sum^{j}_{m_{1},m_{2}=-j}(-1)^{j-m_{2}}\langle j,m_{1},j,m_{2}|j,j,l,m\rangle\bar{\varphi}_{n_{s}km},$ (64)
where
$n=2j+1,\quad m=m_{1}+m_{2},$ (65)
$k=\left\{\begin{aligned} &m_{2}-m_{1},\quad m_{2}-m_{1}>0\ ||\ m_{1}=m_{2}\neq 0,\\ &j,\quad m_{1}=m_{2}=0,\\ &2j-|m_{1}+m_{2}|+m_{1}-m_{2},\quad m_{2}-m_{1}>0,\end{aligned}\right$.
and $\langle j,m_{1},j,m_{2}|j,j,l,m\rangle$ denotes the Clebsch-Gordan coefficients. We refer readers to Ref. [54] for the derivation of the relation.
Appendix C RELATION BETWEEN THE ATOMIC EIGENSTATES AND THE MOLECULAR EIGENSTATES
In this appendix, we demonstrate how to obtain the relation between the atomic eigenstates and the molecular eigenstates in the large separation limit with the prescription described in Sec. V.1. For illustration, we shall consider the two spin-aligned states $\varphi_{211}$, and derive Eq. (31) step by step. We shall refer to the spin directions of the two black holes as $\hat{\bf z}_{1}$ and $\hat{\bf z}_{2}$ and the direction of $\mathbf{r_{2}-r_{1}}$ as $\hat{\bf x}$. Since the two atomic states have the same eigenenergy, the involved molecular states should also have the same eigenenergy. Therefore, we can focus on the spatial part of the wave functions.
At large orbital separation, the wave function is just the superposition of two atomic states
$\psi_{0}=\varphi_{1,211}+\varphi_{2,211}\,.$ (66)
By changing to the corotating frame, it introduces an extra phase,
$\psi_{0}=e^{-i\Omega t}\left(\bar{\varphi}_{1,211}+\bar{\varphi}_{2,211}\right).$ (67)
Note that $\bar{\varphi}_{1,211}$ and $\bar{\varphi}_{2,211}$ are eigenstates of angular momentum in $\hat{\bf z}_{1}$ and $\hat{\bf z}_{2}$, while the molecular eigenstates are eigenstates of angular momentum in $\hat{\bf x}$. We first decompose $\bar{\varphi}_{1,211}$ and $\bar{\varphi}_{2,211}$ into the atomic eigenstates of angular momentum in $\hat{\bf x}$,
[…]
Next, we decompose each atomic eigenstate into atomic eigenstates in parabolic coordinates using Eq. (64),
$\psi_{0}=e^{-i\Omega_{0}t}\sum_{i=1,2}\left(-\frac{1}{2}\bar{\varphi}^{(p)}_{i,20-1}+\frac{1}{2}\bar{\varphi}^{(p)}_{i,200}-\frac{1}{2}\bar{\varphi}^{(p)}_{i,210}+\frac{1}{2}\bar{\varphi}^{(p)}_{i,201}\right)$. (69)
Finally, using the relation (13) derived from node conservation [55], we can express
$\bar{\varphi}^{(p)}_{1,{\rm n_{s}km}}+\bar{\varphi}^{(p)}_{2,{\rm n_{s}km}}=\sqrt{2}\bar{\psi}_{n\ell m}\,.$ (70)
In particular, we have
$\bar{\varphi}^{(p)}_{1,{\rm 20-1}}+\bar{\varphi}^{(p)}_{2,{\rm 20-1}}=\sqrt{2}\bar{\psi}_{21-1}\,$$\quad$ $\bar{\varphi}^{(p)}_{1,{\rm 200}}+\bar{\varphi}^{(p)}_{2,{\rm 200}}=\sqrt{2}\bar{\psi}_{310}\,$, (71)
$\bar{\varphi}^{(p)}_{1,{\rm 210}}+\bar{\varphi}^{(p)}_{2,{\rm 210}}=\sqrt{2}\bar{\psi}_{430}\,$$\quad$ $\bar{\varphi}^{(p)}_{1,{\rm 201}}+\bar{\varphi}^{(p)}_{2,{\rm 201}}=\sqrt{2}\bar{\psi}_{211}\,$,
substituting which into Eq. (69) leads to
$\psi_{0}=\frac{e^{-i\Omega_{0}t}}{\sqrt{2}}\left(\bar{\psi}_{211}-\bar{\psi}_{21-1}+\bar{\psi}_{310}-\bar{\psi}_{430}\right)\,$, (72)
and hence Eq. (31). Equation (LABEL:eq:init2) can be obtained in the same way. One can further express a single atomic state, e.g., $\varphi_{1,211}$ in terms of molecular states, and hence construct the initial condition even if the clouds surrounding the two black holes are in different atomic eigenstates or with nonparallel spin alignment. Nevertheless, this relation is valid only for equal-mass binaries.